# The N-prime graph and the Subgroup Isomorphism Problem

### v1

@@ Abstract.
We introduce a directed graph related to a group $G$, which we call the N-prime graph $\Gamma_{\rm{N}}(G)$ of $G$ and which is a refinement of the classical Gruenberg-Kegel graph. The vertices of $\Gamma_{\rm{N}}(G)$ are the primes $p$ such that $G$ has an element of order $p$, and, for distinct vertices $p$ and $q$, the arc $q\rightarrow p$ is in the graph if and only if $G$ has a subgroup of order $p$ whose normalizer in $G$ has an element of order $q$. Generalizing some known results about the Gruenberg-Kegel graph, we prove that the group $V({\mathbb{Z}}G)$ of the units with augmentation $1$ in the integral group ring ${\mathbb{Z}}G$ has the same N-prime graph as $G$ if $G$ is a finite solvable group, and we reduce to almost simple groups the problem of whether $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))=\Gamma_{\rm{N}}(G)$ holds for any finite group $G$. We also prove that $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))=\Gamma_{\rm{N}}(G)$ if $G$ is almost simple with socle either an alternating group, or ${\rm PSL}_{2}(r^{f})$ with $r$ prime and $f\leq 2$. Finally, for $G$ solvable we obtain some stronger results which give a contribution to the Subgroup Isomorphism Problem. More precisely, we prove that if $V({\mathbb{Z}}G)$ contains a Frobenius subgroup $T$ with kernel of prime order and complement of prime power order, then $G$ contains a subgroup isomorphic to $T$.
@@ 1. Introduction
This paper deals with the following general question concerning finite groups: how much information about a finite group $G$ can be obtained from the ring theoretical properties of its integral group ring ${\mathbb{Z}}G$? The history of this question began when G. Higman proved in his Ph.D. dissertation that if $G$ is abelian, then every group $H$ such that ${\mathbb{Z}}G$ and ${\mathbb{Z}}H$ are isomorphic must be isomorphic to $G$ [15, 16]. A. Whitcomb extended the latter result to the case when $G$ is metabelian [27], and later A. Weiss proved the same property for nilpotent groups [26]. For a while it was believed that this could be true for every finite group but, in 2001, M. Hertweck discovered two non-isomorphic groups with isomorphic integral group rings [10]. Still many evidences show that much information about $G$ is encoded in ${\mathbb{Z}}G$, and mostly in the group $V({\mathbb{Z}}G)$ of units of ${\mathbb{Z}}G$ of augmentation 1. A detailed discussion about the interplay between properties of $G$ and $V({\mathbb{Z}}G)$ can be found in the survey [24].
@@
The First Zassenhaus Conjecture (formulated by H.J. Zassenhaus in the 70’s) is an iconic problem in this research field, which attracted the interest of many authors in the past years; it predicts that every torsion unit of ${\mathbb{Z}}G$ is conjugate in ${\mathbb{Q}}G$ to an element of the form $\pm g$ for a suitable $g\in G$. This was proved in many special cases (see e.g. [1, 5, 13, 26]) but, in 2018, F. Eisele and L. Margolis provided a counterexample [8]. As a related concept, we recall that the Gruenberg-Kegel graph of a group $G$ (also known as the prime graph of $G$) is the simple undirected graph $\Gamma_{\rm{GK}}(G)$ whose vertex set is the set $\pi(G)$ of the prime numbers dividing the order of some torsion element of $G$, and, for distinct vertices $p$ and $q$, the edge qpq-p is in the graph if and only if $G$ has an element of order $pq$. Now, if the First Zassenhaus Conjecture holds for a given group $G$, then clearly $V({\mathbb{Z}}G)$ and $G$ have the same Gruenberg-Kegel graph. In view of this, W. Kimmerle proposed in [20] the so-called Prime Graph Question: is it true that, for every finite group $G$, the graphs $\Gamma_{\rm{GK}}(V({\mathbb{Z}}G))$ and $\Gamma_{\rm{GK}}(G)$ are equal? The Prime Graph Question is usually abbreviated as (PQ), referring to the common use of $p$ and $q$ for two different primes. It can be interpreted as a weaker form of the First Zassenhaus Conjecture, and there is no known counterexample for it. Kimmerle proved in [20] that the answer is affirmative if $G$ is a solvable group, and in fact the problem has been reduced to almost simple groups by Kimmerle and A. Konovalov in [21]; since then, several almost simple groups have been investigated from this point of view (see [24, Theorem 2.6] for a wide list of results on the Prime Graph Question for almost simple groups).
@@
In this paper we introduce a different graph related to a group $G$, which we call the N-prime graph $\Gamma_{\rm{N}}(G)$ of $G$. We define it as the simple directed graph whose vertex set is $\pi(G)$ and, for distinct vertices $p$ and $q$, the arc $q\rightarrow p$ is in the graph if and only if there exists an element $x$ in $G$ of order $p$ such that $\left\langle x\right\rangle$ is normalized by an element $y$ in $G$ of order $q$. As observed in Remark˜3.1, the Gruenberg-Kegel graph of $G$ can be obtained from $\Gamma_{\rm{N}}(G)$ replacing every double arrow $q\leftrightarrows p$ by the edge qpq-p and deleting all single arrows; on the other hand, the converse is not true because it is easy to find examples of groups having the same Gruenberg-Kegel graph but different N-prime graphs. So, this graph in principle encodes more detailed information about the structure of the group $G$, and we can consider the following strengthening of the Prime Graph Question.
@@
In view of ˜A, we explore the validity of (NPQ) for some families of almost simple groups. In particular, we prove that (NPQ) holds for rational groups, and for almost simple groups whose socle is an alternating group or a group of the form ${\rm{PSL}}_{2}(r^{f})$ where $r$ is a prime and $f\in\{1,2\}$ (see Theorem˜3.5, Theorem˜4.1 and Theorem˜4.3, respectively). As a consequence, we obtain the following.
@@
Another interesting problem in the present context, somewhat “intermediate" between the First Zassenhaus Conjecture and the Prime Graph Question, is the Spectrum Problem: is it true that, whenever $V({\mathbb{Z}}G)$ has an element of a given order $n$, then $G$ has an element of order $n$ as well? It has been proved by Hertweck in [11] that the answer is affirmative if the group $G$ is solvable, and there is no known counterexample. By the Cohn-Livingstone Theorem [6], we know that if $V({\mathbb{Z}}G)$ has a cyclic subgroup $T$ of prime-power order, then also $G$ has a subgroup isomorphic to $T$. Note that the Spectrum Problem (respectively, the Prime Graph Question) can be restated as follows: is it true that if $V({\mathbb{Z}}G)$ has a cyclic subgroup $C_{n}$ of order $n$ (respectively, of order $pq$), then so does $G$? So these questions ask for the isomorphism types of cyclic subgroups of $G$ and $V({\mathbb{Z}}G)$. On the other hand, (NPQ) deals with subgroups of type $C_{p}\rtimes C_{q}$. We can frame all these questions in the following general context.
@@
A caution must be taken into account to consider (NPQ) as part of the (SIP). Let $T$ be the Frobenius group $C_{p}\rtimes C_{q}$. Suppose that (NPQ) holds for $G$ and $V({\mathbb{Z}}G)$ has a subgroup isomorphic to $T$. If, moreover, $G$ does not have elements of order $pq$, then $G$ has a subgroup isomorphic to $T$. However, if $G$ has elements of order $pq$, then the existence of a subgroup of $G$ isomorphic to $T$ is not granted by (NPQ).
@@ 2. Preliminary notation
(see [25, (41.1), page 237]), which implies that $\varepsilon_{b}(u)=\varepsilon_{b}(v)$ if $u$ and $v$ are conjugate in $RG$.
@@
3. The N-Prime Graph Question: ˜A
@@
If $C$ centralizes $P$, then we can easily construct an element of order $pq$ of $G$. Otherwise, Lemma˜3.2 yields the existence of a subgroup $U$ of $P$ such that $|U|=p$ and $C\subseteq{\bf N}_{G}(U)$. Now the subgroup $UC$ yields the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(G)$, and the proof is complete.
@@
We are ready to start our discussion about the N-Prime Graph Question, that was presented in the introduction. Note that, by the Cohn-Livingstone Theorem [6], $\Gamma_{\rm{N}}(G)$ and $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ have the same set of vertices. Moreover, it is clear that $\Gamma_{\rm{N}}(G)$ is a subgraph of $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$. So, the N-Prime Graph Question amounts to understand if, for $p,q$ distinct primes in $\pi(G)$, the existence of the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ implies the existence of the same arc in $\Gamma_{\rm{N}}(G)$.
@@
The following result, which collects together ˜A and ˜B, shows that the N-Prime Graph Question can be reduced to almost simple groups. This extends the corresponding result by Kimmerle and Konovalov for the Prime Graph Question.
@@
As observed in the paragraph after Remark˜3.4, we only have to show that, for $p$ and $q$ distinct primes in $\pi(G)$, the existence of the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ implies the existence of the same arc in $\Gamma_{\rm{N}}(G)$. For a proof by contradiction, we will assume that $G$ is a minimal counterexample to this fact with respect to the primes $p$ and $q$. Note that, by our assumptions and by Remark˜3.1, if $X$ is any almost simple homomorphic image of $G$ then also the Gruenberg-Kegel graphs $\Gamma_{\rm{GK}}(X)$ and $\Gamma_{\rm{GK}}(V({\mathbb{Z}}X))$ coincide; therefore, by [21], we have $\Gamma_{\rm{GK}}(G)=\Gamma_{\rm{GK}}(V({\mathbb{Z}}G))$, and the presence of both arcs $q\leftrightarrows p$ in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ would imply the presence of the edge qpq-p in $\Gamma_{\rm{GK}}(G)$. As a consequence both arcs $q\leftrightarrows p$ would be in $\Gamma_{\rm{N}}(G)$, against the fact that $G$ is a counterexample. So $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ has the arc $q\rightarrow p$ but not the arc $p\rightarrow q$; in particular, $V({\mathbb{Z}}G)$ has a subgroup $T$ which is a Frobenius group of order $pq$, and $q$ is a divisor of p1p-1.
Let $M$ be a minimal normal subgroup of $G$, and observe that the factor group $G/M$ satisfies our assumptions. Denoting by $\varphi:V({\mathbb{Z}}G)\to V({\mathbb{Z}}(G/M))$ the natural homomorphism, if $p$ does not divide $|M|$, then we get $T\cap{\rm ker}(\varphi)=1$ (see [25, Lemma 7.5]) and hence $\varphi(T)\cong T$. It follows that $q\rightarrow p$ is an arc of $\Gamma_{\rm{N}}(V({\mathbb{Z}}(G/M)))$ and, by our minimality assumption, it is also an arc of $\Gamma_{\rm{N}}(G/M)$. But then $q\rightarrow p$ is an arc of $\Gamma_{\rm{N}}(G)$ by Proposition˜3.3, a contradiction. Therefore, no minimal normal subgroup of $G$ can be a $p^{\prime}$-group.
The discussion in the paragraph above implies that the Fitting subgroup ${\bf F}(G)$ of $G$ is a $p$-group, and we claim that in fact ${\bf F}(G)$ must be trivial. As in the proof of Proposition˜3.3, consider a subgroup $C$ of $G$ having order $q$ and assume ${\bf F}(G)\neq 1$. If $C$ centralizes ${\bf F}(G)$ then we can construct an element of order $pq$ of $G$, not our case. Otherwise, we can use Lemma˜3.2 and find a subgroup $U$ of ${\bf F}(G)$ such that $|U|=p$ and $C\subseteq{\bf N}_{G}(U)$. Now the subgroup $UC$ yields the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(G)$, again a contradiction.
@@
As another remark, given a minimal normal subgroup $M\cong S_{1}\times\cdots\times S_{t}$ of $G$, for any fixed $i\in\{1,\ldots,t\}$ we claim that $q$ does not divide $|{\bf N}_{G}(S_{i})|$. Assuming the contrary, we could take an element $y$ of order $q$ in ${\bf N}_{G}(S_{i})$. Recalling that $S_{i}$ is a $q^{\prime}$-group having an order divisible by $p$, we could find a (nontrivial) Sylow $p$-subgroup $P$ of $S_{i}$ such that $y$ normalizes $P$ (see [22, 8.2.3(a)]), and the same argument as in the third paragraph of this proof leads to a contradiction.
@@
The following corollary can be immediately derived from the theorem above and from Theorem˜3.5.
@@
4. The N-Prime Graph Question: ˜C
In this section we prove that NPQ holds for every almost simple group whose socle is isomorphic to an alternating group, or to ${\rm{PSL}}_{2}(r^{f})$ where $r$ is a prime and $f\in\{1,2\}$ (Theorem˜4.1 and Theorem˜4.3, respectively); this, together with Theorem˜3.6 and Corollary˜3.7, will conclude the proof of ˜C of the introduction.

### v2

@@ Abstract.
We introduce a directed graph related to a group $G$, which we call the N-prime graph $\Gamma_{\rm{N}}(G)$ of $G$ and which is a refinement of the classical Gruenberg-Kegel graph. The vertices of $\Gamma_{\rm{N}}(G)$ are the primes $p$ such that $G$ has an element of order $p$, and, for distinct vertices $p$ and $q$, the arc $q\rightarrow p$ is in the graph if and only if $G$ has a subgroup of order $p$ whose normalizer in $G$ has an element of order $q$. Generalizing some known results about the Gruenberg-Kegel graph, we prove that the group $V({\mathbb{Z}}G)$ of the units with augmentation $1$ in the integral group ring ${\mathbb{Z}}G$ has the same N-prime graph as $G$ if $G$ is a finite solvable group, and we reduce to almost simple groups the problem of whether $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))=\Gamma_{\rm{N}}(G)$ holds for any finite group $G$. We also prove that $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))=\Gamma_{\rm{N}}(G)$ if $G$ is almost simple with socle either an alternating group, or ${\rm PSL}_{2}(r^{f})$ with $r$ prime and $f\leq 2$. Finally, for $G$ solvable we obtain some stronger results which give a contribution to the Subgroup Isomorphism Problem. More precisely, we prove that if $V({\mathbb{Z}}G)$ contains a Frobenius subgroup $T$ with kernel of prime order and complement of prime-power order, then $G$ contains a subgroup isomorphic to $T$.
@@ 1. Introduction
This paper deals with the following general question concerning finite groups: how much information about a finite group $G$ can be obtained from the ring theoretical properties of its integral group ring ${\mathbb{Z}}G$? The history of this question began when G. Higman proved in his Ph.D. dissertation that if $G$ is abelian, then every group $H$ such that ${\mathbb{Z}}G$ and ${\mathbb{Z}}H$ are isomorphic must be isomorphic to $G$ [16, 17]. A. Whitcomb extended the latter result to the case when $G$ is metabelian [29], and later A. Weiss proved the same property for nilpotent groups [28]. For a while it was believed that this could be true for every finite group but, in 2001, M. Hertweck discovered two non-isomorphic groups with isomorphic integral group rings [11]. Nevertheless, there is still much evidence that significant information about $G$ is encoded in ${\mathbb{Z}}G$, and mostly in the group $V({\mathbb{Z}}G)$ of units of ${\mathbb{Z}}G$ of augmentation 1. A detailed discussion about the interplay between properties of $G$ and $V({\mathbb{Z}}G)$ can be found in the survey [26] and in [20].
@@
The First Zassenhaus Conjecture (formulated by H.J. Zassenhaus in 1974) is an iconic problem in this research field, which attracted the interest of many authors in the past years; it predicts that every torsion unit of ${\mathbb{Z}}G$ is conjugate in ${\mathbb{Q}}G$ to an element of the form $\pm g$ for a suitable $g\in G$. This was proved in many special cases (see e.g. [1, 5, 14, 28]) but, in 2018, F. Eisele and L. Margolis provided a counterexample [9]. As a related concept, we recall that the Gruenberg-Kegel graph of a group $G$ (also known as the prime graph of $G$) is the simple undirected graph $\Gamma_{\rm{GK}}(G)$ whose vertex set is the set $\pi(G)$ of the prime numbers dividing the order of some torsion element of $G$, and, for distinct vertices $p$ and $q$, the edge qpq-p is in the graph if and only if $G$ has an element of order $pq$. Now, if the First Zassenhaus Conjecture holds for a given group $G$, then clearly $V({\mathbb{Z}}G)$ and $G$ have the same Gruenberg-Kegel graph. In view of this, W. Kimmerle proposed in [22] the so-called Prime Graph Question: is it true that, for every finite group $G$, the graphs $\Gamma_{\rm{GK}}(V({\mathbb{Z}}G))$ and $\Gamma_{\rm{GK}}(G)$ are equal? The Prime Graph Question is usually abbreviated as (PQ), referring to the common use of $p$ and $q$ for two different primes. It can be interpreted as a weaker form of the First Zassenhaus Conjecture, and there is no known counterexample for it. Kimmerle proved in [22] that the answer is affirmative if $G$ is a solvable group, and in fact the problem has been reduced to almost simple groups by Kimmerle and A. Konovalov in [23]; since then, several almost simple groups have been investigated from this point of view (see [26, Theorem 2.6] for a wide list of results on the Prime Graph Question for almost simple groups; see also [8]).
@@
In this paper we introduce a different graph related to a group $G$, which we call the N-prime graph $\Gamma_{\rm{N}}(G)$ of $G$. We define it as the simple directed graph whose vertex set is $\pi(G)$ and, for distinct vertices $p$ and $q$, the arc $q\rightarrow p$ is in the graph if and only if there exists an element $x$ in $G$ of order $p$ such that $\left\langle x\right\rangle$ is normalized by an element $y$ in $G$ of order $q$. As observed in Remark 3.1, the Gruenberg-Kegel graph of $G$ can be obtained from $\Gamma_{\rm{N}}(G)$ replacing every double arrow $q\leftrightarrows p$ by the edge qpq-p and deleting all single arrows; on the other hand, the converse is not true because it is easy to find examples of groups having the same Gruenberg-Kegel graph but different N-prime graphs. So, this graph in principle encodes more detailed information about the structure of the group $G$, and we can consider the following strengthening of the Prime Graph Question.
@@
In view of Theorem A, we explore the validity of (NPQ) for some families of almost simple groups. In particular, we prove that (NPQ) holds for rational groups, and for almost simple groups whose socle is an alternating group or a group of the form ${\rm{PSL}}_{2}(r^{f})$ where $r$ is a prime and $f\in\{1,2\}$ (see Theorem 3.5, Theorem 4.1 and Theorem 4.3, respectively). As a consequence, we obtain the following.
@@
Another interesting problem in the present context, somewhat “intermediate" between the First Zassenhaus Conjecture and the Prime Graph Question, is the Spectrum Problem: is it true that, whenever $V({\mathbb{Z}}G)$ has an element of a given order $n$, then $G$ has an element of order $n$ as well? It has been proved by Hertweck in [12] that the answer is affirmative if the group $G$ is solvable, and there is no known counterexample. By Corollary 4.1 in [6], we know that if $V({\mathbb{Z}}G)$ has a cyclic subgroup $T$ of prime-power order, then also $G$ has a subgroup isomorphic to $T$. Note that the Spectrum Problem (respectively, the Prime Graph Question) can be restated as follows: is it true that if $V({\mathbb{Z}}G)$ has a cyclic subgroup $C_{n}$ of order $n$ (respectively, of order $pq$), then so does $G$? So these questions ask for the isomorphism types of cyclic subgroups of $G$ and $V({\mathbb{Z}}G)$. On the other hand, (NPQ) deals with subgroups of type $C_{p}\rtimes C_{q}$. We can frame all these questions in the following general context.
@@
Care should be taken in considering (NPQ) as part of the (SIP). Let $T$ be the Frobenius group $C_{p}\rtimes C_{q}$. Suppose that (NPQ) holds for $G$ and $V({\mathbb{Z}}G)$ has a subgroup isomorphic to $T$. If, moreover, $G$ does not have elements of order $pq$, then $G$ has a subgroup isomorphic to $T$. However, if $G$ has elements of order $pq$, then the existence of a subgroup of $G$ isomorphic to $T$ is not granted by (NPQ).
@@ 2. Preliminary notation
(see [27, (41.1), page 237]), which implies that $\varepsilon_{b}(u)=\varepsilon_{b}(v)$ if $u$ and $v$ are conjugate in $RG$.
@@
3. The N-Prime Graph Question: Theorem A
@@
If $C$ centralizes $P$, then we can easily construct an element of order $pq$ of $G$. Otherwise, Lemma 3.2 yields the existence of a subgroup $U$ of $P$ such that $|U|=p$ and $C\subseteq{\bf N}_{G}(U)$. Now the subgroup $UC$ yields the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(G)$, and the proof is complete.
@@
We are ready to start our discussion about the N-Prime Graph Question, that was presented in the introduction. Note that, by [6, Corollary 4.1], $\Gamma_{\rm{N}}(G)$ and $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ have the same set of vertices. Moreover, it is clear that $\Gamma_{\rm{N}}(G)$ is a subgraph of $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$. So, the N-Prime Graph Question amounts to understanding if, for $p,q$ distinct primes in $\pi(G)$, the existence of the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ implies the existence of the same arc in $\Gamma_{\rm{N}}(G)$.
@@
The following result, which collects together Theorem A and Corollary B, shows that the N-Prime Graph Question can be reduced to almost simple groups. This extends the corresponding result by Kimmerle and Konovalov for the Prime Graph Question.
@@
As observed in the paragraph after Remark 3.4, we only have to show that, for $p$ and $q$ distinct primes in $\pi(G)$, the existence of the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ implies the existence of the same arc in $\Gamma_{\rm{N}}(G)$. For a proof by contradiction, we will assume that $G$ is a minimal counterexample to this fact with respect to the primes $p$ and $q$. Note that, by our assumptions and by Remark 3.1, if $X$ is any almost simple homomorphic image of $G$ then also the Gruenberg-Kegel graphs $\Gamma_{\rm{GK}}(X)$ and $\Gamma_{\rm{GK}}(V({\mathbb{Z}}X))$ coincide; therefore, by [23], we have $\Gamma_{\rm{GK}}(G)=\Gamma_{\rm{GK}}(V({\mathbb{Z}}G))$, and the presence of both arcs $q\leftrightarrows p$ in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ would imply the presence of the edge qpq-p in $\Gamma_{\rm{GK}}(G)$. As a consequence both arcs $q\leftrightarrows p$ would be in $\Gamma_{\rm{N}}(G)$, against the fact that $G$ is a counterexample. So $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ has the arc $q\rightarrow p$ but not the arc $p\rightarrow q$; in particular, $V({\mathbb{Z}}G)$ has a subgroup $T$ which is a Frobenius group of order $pq$, and $q$ is a divisor of p1p-1.
Let $M$ be a minimal normal subgroup of $G$, and observe that the factor group $G/M$ satisfies our assumptions. Denoting by $\varphi:V({\mathbb{Z}}G)\to V({\mathbb{Z}}(G/M))$ the natural homomorphism, if $p$ does not divide $|M|$, then we get $T\cap{\rm ker}(\varphi)=1$ (see [27, Lemma 7.5]) and hence $\varphi(T)\cong T$. It follows that $q\rightarrow p$ is an arc of $\Gamma_{\rm{N}}(V({\mathbb{Z}}(G/M)))$ and, by our minimality assumption, it is also an arc of $\Gamma_{\rm{N}}(G/M)$. But then $q\rightarrow p$ is an arc of $\Gamma_{\rm{N}}(G)$ by Proposition 3.3, a contradiction. Therefore, no minimal normal subgroup of $G$ can be a $p^{\prime}$-group.
The discussion in the paragraph above implies that the Fitting subgroup ${\bf F}(G)$ of $G$ is a $p$-group, and we claim that in fact ${\bf F}(G)$ must be trivial. As in the proof of Proposition 3.3, consider a subgroup $C$ of $G$ having order $q$ and assume ${\bf F}(G)\neq 1$. If $C$ centralizes ${\bf F}(G)$ then we can construct an element of order $pq$ of $G$, not our case. Otherwise, we can use Lemma 3.2 and find a subgroup $U$ of ${\bf F}(G)$ such that $|U|=p$ and $C\subseteq{\bf N}_{G}(U)$. Now the subgroup $UC$ yields the arc $q\rightarrow p$ in $\Gamma_{\rm{N}}(G)$, again a contradiction.
@@
As another remark, given a minimal normal subgroup $M\cong S_{1}\times\cdots\times S_{t}$ of $G$, for any fixed $i\in\{1,\ldots,t\}$ we claim that $q$ does not divide $|{\bf N}_{G}(S_{i})|$. Assuming the contrary, we could take an element $y$ of order $q$ in ${\bf N}_{G}(S_{i})$. Recalling that $S_{i}$ is a $q^{\prime}$-group having an order divisible by $p$, we could find a (nontrivial) Sylow $p$-subgroup $P$ of $S_{i}$ such that $y$ normalizes $P$ (see [24, 8.2.3(a)]), and the same argument as in the third paragraph of this proof leads to a contradiction.
@@
The following corollary can be immediately derived from the theorem above and from Theorem 3.5.
@@
4. The N-Prime Graph Question: Theorem C
In this section we prove that (NPQ) holds for every almost simple group whose socle is isomorphic to an alternating group, or to ${\rm{PSL}}_{2}(r^{f})$ where $r$ is a prime and $f\in\{1,2\}$ (Theorem 4.1 and Theorem 4.3, respectively); this, together with Theorem 3.6 and Corollary 3.7, will conclude the proof of Theorem C of the introduction.